OpenStudy (anonymous):

the length of rectangular garden 6 meters, is 21 meters of fencing are required to fence the garden .find its width

OpenStudy (issy14):

 In order to solve this, you have to think about what 21m represent, does it represent the area of the perimeter? By utilizing the area formula of a rectangle you can find out the answer. a= wl ( substitute: 21=w6) arrange the formula: 21/6 = w 3.4=w. Verify it : a=6*3.5= 21

OpenStudy (anonymous):

thanks alot

OpenStudy (issy14):

No problem

OpenStudy (anonymous):

one more question

OpenStudy (issy14):

yup

OpenStudy (mrnood):

@arslan @Issy14 The question states that the PERIMETER is 21m (21m of fence required to surround the garden - we have to assume all 4 sides are fenced...) The perimeter is = 2L + 2W S0 2L+2W = 21 You know L so you can solve for W